# CoLA: Compute-Efficient Pre-Training of LLMs via Low-Rank Activation

### v1

@@ Abstract
Large language models (LLMs) are revolutionizing many science and engineering fields. However, their huge model sizes impose extremely demanding needs of computational resources in the pre-training stage. Although low-rank factorizations can reduce model parameters, their direct application in LLM pre-training often lead to non-negligible performance loss. To address this fundamental challenge, we introduce CoLA and its memory-efficient implementation, CoLA-M. We leverage the low-rank structure observed widely in model activations, enforcing non-linear transformations between factorized weight matrices to reduce model size, boost model capacity and training efficiency. Experiments on LLaMA models with 60 million to 7 billion parameters show that CoLA reduces the computing cost by $\bf 2\boldsymbol{\times}$ and improves training throughput by $\bf 1.86\boldsymbol{\times}$ while maintaining full-rank level performance. CoLA-M further squeezes memory cost without sacrificing throughput, offering a pre-training approach with collectively superior parameter, computing, and memory efficiency. The LLMs produced are also $\bf 2\boldsymbol{\times}$ smaller, enabling faster inference with lower memory cost on resource-constrained platforms.  [^1] Code available here.
@@ 1 Introduction
Figure 1: A joint comparison of validation perplexity, estimated compute FLOPs (per block, step) and model size between various pre-training methods on a LLaMA-1B model with a token batch size of 256. Among them, our proposed CoLA is the only one that reduces both compute FLOPs and model size while demonstrating on par validation perplexity with full-rank training.
@@
Motivation: At the core of increasing resource utilization and cost is the simple practice of scaling up full-size linear layers in decoder-only architectures, which has proven to be a viable and straightforward strategy. Thus, to break free from this unsustainable trend, it is imperative to improve architecture efficiency. This has been widely studied in the deep learning community, involving different levels of factorization of weight matrices: from simple matrix factorizations, i.e., a singular value decomposition (SVD), to higher-order tensor factorizations such as Canonical Polyadic, Tucker, and Tensor-Train (TT) format. Extensive studies have shown that such factorizations can effectively reduce the total number of parameters needed to achieve similar performance in numerous domains Sainath et al. (2013); Jaderberg et al. (2014); Lebedev et al. (2014); Novikov et al. (2015); Tjandra et al. (2017); Dao et al. (2021); Sui et al. (2024); Yang et al. (2024); Zhang et al. (2024), especially when neural networks are overparameterized.
@@
Limitations of state-of-art: The techniques mentioned above have been applied only to a limited degree to pre-training tasks, and their findings suggest that the pure low-rank or sparse structure often downgrades model performance Khodak et al. (2021); Kamalakara et al. (2022); Chekalina et al. (2023); Zhao et al. (2024); Hu et al. (2024); Mozaffari et al. (2024). This has pivoted most recent work of efficient pre-training into two directions: 1) Accumulating multiple low-rank updates Huh et al. (2024); Lialin et al. (2023); 2) Enforcing low-rank structures in gradients rather than parameters Zhao et al. (2024); Chen et al. (2024); Huang et al. ; Liao et al. (2024); Hao et al. (2024); Zhu et al. (2024). Both approaches have their limitations. 1) The accumulation of low-rank updates requires instantiating a full-rank matrix and a deeply customized training strategy that periodically merges and restarts the low-rank components. This creates computing overhead in practice and can only achieve (if only) marginal computing and memory reduction. 2) Enforcing low-rank gradients reduces only the optimizer memory and adds additional computation that downgrades training throughput. Furthermore, the memory saving caused by gradient compression becomes negligible as the training batch size increases, as activations dominate the total memory cost. A recent paper called SLTrain Han et al. (2024) revisited the notion of parameter efficiency in foundation model pre-training, by having both low-rank factors and an unstructured sparse matrix. SLTrain effectively reduces the total number of parameters without significantly hurting model performance. However, it still introduces computing overhead on top of full-rank training due to the necessary reconstruction of low-rank factors. We note that none of the above works has achieved superior efficiency of parameter, computing, and memory simultaneously without performance drop in both training and inference for foundation model pre-training.
@@
Contributions: In this paper, we propose CoLA: Compute-Efficient Pre-Training of LLMs via Low-rank Activation, and its memory efficient implementation CoLA-M, to achieve all the desirable properties mentioned above. We summarize our contributions as follows:
  • We propose CoLA, a novel architecture that enforces explicit low-rank activations by injecting non-linear operations between factorized weight matrices. CoLA can greatly reduce the computing FLOPS while maintaining the performance of full-rank pre-training.
@@
  • We extensively pre-train LLaMA with 60M up to 7B parameters. CoLA reduces model size and computing FLOPs by $\bf 2\boldsymbol{\times}$, while maintaining on-par performance to its full-rank counterpart. At the system level, CoLA improves $\bf 1.86\boldsymbol{\times}$ training and $\bf 1.64\boldsymbol{\times}$ inference throughput. CoLA-M reduces total pre-training memory by $\bf 2/3$, while still manages to improve $\bf 1.3\boldsymbol{\times}$ training throughput over full-rank baselines.
A high-level comparison of CoLA/CoLA-M with main baselines is provided in Table 1.
@@ 2 Related Work
Model Compression. Recent research on efficient LLM pre-training primarily focuses on memory savings. To out best knowledge, SLTrain Han et al. (2024) is the first method that reduces both trainable parameters and total parameters in LLM pre-training, without significantly hurting model performance. This also reduces memory usage for model, gradients, and optimizer states (see its smaller circle in Fig. 1). However, the existence of its unstructured sparse matrix $\mathbf{S}$ requires reconstructing $\tilde{\mathbf{W}}=\mathbf{BA}+\mathbf{S}$, otherwise it will incur dense-sparse multiplications that are still memory costly (Fig. 3c). This causes additional computing than the full-rank baseline. LoRA/ReLoRA Hu et al. (2021); Lialin et al. (2023) reduces trainable parameters by freezing a full-rank $\mathbf{W}_{0}$ and training (at least in a later stage) only low-rank factors, potentially reducing memory needs. Yet, any compute savings are limited because the forward pass yields a larger compute than its full-rank counterpart, especially when the rank must stay relatively large in pre-training tasks. CoMERA Yang et al. (2024) achieves higher model compression and FLOP reduction, but its low-rank tensor operations are GPU unfriendly. Similar to matrix-compressed approaches, CoMERA cannot avoid a performance drop either. Some works investigate pure structured sparsity or combined with low-rank factors Hu et al. (2024); Mozaffari et al. (2024) to achieve speed up, but still show a significant performance drop during the pre-training stage.
@@
Gradient Compression. GaLore Zhao et al. (2024) reduces memory by projecting gradients into a low-rank space, shrinking optimizer states below the typical 2×2$\times$2 × AdamW overhead Loshchilov (2017). However, it increases computation by adding up/down projections on top of already compute-heavy full-rank training. As shown in Fig. 1, its estimated FLOPs surpass full-rank training on the LLaMA-1B scale. Follow-up work Chen et al. (2024); Huang et al. ; Liao et al. (2024); Hao et al. (2024); Zhu et al. (2024) further explores low-rank gradient projection. While these methods are promising, they are mostly orthogonal to our focus. Crucially, these methods are still computing lower-bounded by the full-rank baseline. Our goal instead is to reduce computing cost to a fraction of full-rank training, thus lowering the demand of computing resources in LLM pre-training.
@@
This paper presents an alternative approach that explores the low-rank property in model activations from an architectural perspective. This is conceptually different from the above model compression methods despite the similarity in their formulations. Our approach is mostly orthogonal with gradient compression techniques, meaning that they can be combined to further boost efficiency.
@@ 3 CoLA for Efficient LLM Pre-Training
Figure 3: Comparison between CoLA (ours) and other efficient pre-training frameworks. a) LoRA/ReLoRA Lialin et al. (2023) maintains a full-rank frozen weight; b) GaLore Zhao et al. (2024) only reduces optimizer states by down and up projecting gradients; c) SLTrain Han et al. (2024) requires necessary reconstruction of the low-rank and sparse matrices; d) CoLA (ours) is a pure low-rank method involves only rank $r$ weight matrices.
@@ 3.1 A Motivating Example
Many previous works have observed the low-rank structure of model activations in deep neural networks Cui et al. (2020); Huh et al. (2021). We also observe this phenomenon in LLMs, i.e. the effective rank of the activations is much smaller than their original dimensionality. To quantify this, we define the effective rank $r(\alpha)$ of activation as the minimal number of singular values needed to preserve an $\alpha$-fraction of the total spectral energy. Formally:
@@
3.2 Low Rank Weights + Activations
Figure 4: A decoder block in CoLA with LLaMA-like architecture (layer norms, rotary positional embeddings are omitted for simplicity). Every linear layer is replaced by low-rank factorization with a non-linear function in between. Modules painted in sketch are the re-computations during the backward step of CoLA-M (a memory efficient implementation of CoLA).
Motivated by the observed low-rank nature of LLM activations, we propose to enforce explicit low-rank activation via injecting non-linearity between factorized weight matrices.
@@ 3.2 Low Rank Weights + Activations
[…]
$\mathbf{h^{\prime}}=\mathbf{B}\,\sigma(\mathbf{A}\mathbf{x}).$ (3)
During the backward step, we simply apply the chain rule to compute gradients w.r.t each of the low rank factors as
$\nabla_{\mathbf{B}}=\nabla_{\mathbf{h}^{\prime}}\mathbf{z}^{T},\nabla_{\mathbf{z}}=\mathbf{B}^{T}\nabla_{\mathbf{h}^{\prime}},\nabla_{\mathbf{o}}=\nabla_{\mathbf{z}}\odot\sigma^{\prime}(\mathbf{o}),$ (4)
$\nabla_{\mathbf{A}}=\nabla_{\mathbf{o}}\mathbf{x}^{T},\nabla_{\mathbf{x}}=\mathbf{A}^{T}\nabla_{\mathbf{o}},$
where $\mathbf{o}=\mathbf{Ax},\mathbf{z}=\sigma(\mathbf{o})$, $\odot$ denote the element-wise product. We empirically find that keeping the original nonlinearity on top of Eq. (3) does not harm the performance, nor necessarily brings benefits. However, applying Eq. (3) to all linear layers regardless of whether being followed by nonlinearity is crucial to boost model performance. We refer more details to the ablation study in Appendix 11.
Fig. 4 shows the architecture of each transformer block when adopting CoLA into the LLaMA architecture. We highlight the fact that only the original linear layers and (if any) their follow-up non-linear transformation are modified to the CoLA formulation. Other computations such as the scaled-dot product of the self-attention, as well as residual connections and the element-wise product of LLaMA’s feed-forward layers, remain unchanged.
@@ 3.3 Computing Efficiency
We summarize the computational costs of these methods in Table 3 and observe that the costs of SLTrain and GaLore are lower bounded by full-rank training, while (Re)LoRA is lower bounded by CoLA when choosing the same rank. In contrast, CoLA reduces the computation from full-rank training when $r<0.62d$, assuming $d_{\text{ff}}\approx 2.5d$ in LLaMA-like architecture. The default rank choice is set to $r=\frac{1}{4}d$, leading to a reduction in compute to about half of the full-rank training. We refer all details of compute analysis to Appendix B.
@@ 4 CoLA-M: A Memory-Efficient Implementation
Although CoLA has more intermediate results from each low-rank projection and the following non-linear function (as shown in Fig. 4), we can choose strategically which ones to save in order to balance re-computations with memory overhead. In this section, we design and develop CoLA-M, a memory-efficient implementation to leverage CoLA’s structural advantage to achieve superior memory saving without sacrificing throughput.
@@ 4.1 Memory Breakdown in Pre-Training
Figure 6: Memory breakdown of pre-training LLaMA-1B on single GPU using different pre-training methods. Activation is at dominance for most methods.
We assume a common notion that training modern transformers with Adam (or AdamW) involves four key memory components Zhao et al. (2024); Han et al. (2024): model parameters, gradients, optimizer states, and activations (intermediate forward-pass results). Gradients consume 1×1$\times$1 × model size, Adam consumes 2×2$\times$2 ×, and activations typically consume 14×1$\sim$ 4$\times$1 ∼ 4 ×, depending on the batch size.
We focus on the scenario where the memory cost determined by the model size is not in the extreme limit of the GPU. We argue that this is rather realistic, since the model size and the minimum required tokens should scale up simultaneously during pre-training Kaplan et al. (2020); Hoffmann et al. (2022); Krajewski et al. (2024); Kumar et al. (2024). A tiny batch size on a single GPU would be impractical. Therefore, we analyze memory usage on a 40-GB A100 or a 94-GB H100 GPU with a fairly large sequence batch size. Fig. 5 shows that activations dominate memory usage in this setup. Although our method overall consumes less memory than full-rank training (largely due to parameter efficiency), vanilla CoLA allocates more memory to activations. However, we argue that our unique architecture can significantly enhance existing memory-saving techniques and consume only a fraction of the original cost.

### v2

@@ Abstract
The full-size MLPs and the projection layers in attention introduce tremendous model sizes of large language models (LLMs), imposing extremely demanding needs of computational resources in the pre-training stage. However, we empirically observe that the activations of pre-trained LLMs exhibit low-rank property. Motivated by such observations, we propose CoLA and its memory-efficient implementation, CoLA-M, to replace these full-size layers with compute-efficient auto-encoders that naturally enforce low-rank activations throughout training. This fundamental architectural change eliminates the activation redundancy and significantly boosts model capacity and training efficiency. Experiments on LLaMA models with 60 million to 7 billion parameters show that CoLA reduces the computing cost by $\bf 2\boldsymbol{\times}$ and improves training throughput by $\bf 1.86\boldsymbol{\times}$ while maintaining full-rank level performance. CoLA-M further squeezes memory cost without sacrificing throughput, offering a pre-training approach with collectively superior parameter, computing, and memory efficiency. The LLMs produced are also $\bf 2\boldsymbol{\times}$ smaller, enabling faster inference with lower memory cost on resource-constrained platforms.  [^1] Code available here.
@@ 1 Introduction
Figure 1: Comparison between various pre-training methods on a LLaMA-1B model with a token batch size of 256. Among them, CoLA is the only one that reduces both compute FLOPs and model size while demonstrating on par validation perplexity with full-rank training.
@@
Motivation: At the core of increasing resource utilization and cost is the simple practice of scaling up full-size linear layers in decoder-only architectures, which has proven to be a viable and straightforward strategy. Thus, to break free from this unsustainable trend, it is imperative to improve architecture efficiency. This has been widely studied in the deep learning community, involving different levels of factorization of weight matrices: from simple matrix factorizations, i.e., a singular value decomposition (SVD), to higher-order tensor factorizations. Extensive studies have shown that such factorizations can effectively reduce the total number of parameters needed to achieve similar performance in numerous domains Jaderberg et al. (2014); Lebedev et al. (2014); Novikov et al. (2015); Tjandra et al. (2017); Dao et al. (2021); Sui et al. (2024); Yang et al. (2024); Zhang et al. (2024), especially when neural networks are overparameterized.
@@
Limitations of state-of-art: The techniques mentioned above have been applied only to a limited degree to pre-training tasks, and their findings suggest that the pure low-rank or sparse structure often downgrades model performance Khodak et al. (2021); Kamalakara et al. (2022); Chekalina et al. (2023); Zhao et al. (2024); Hu et al. (2024); Mozaffari et al. (2024). This has pivoted most recent work of efficient pre-training into two directions: 1) Accumulating multiple low-rank updates Huh et al. (2024); Lialin et al. (2023); Loeschcke et al. (2024); 2) Enforcing low-rank structures in gradients rather than parameters Zhao et al. (2024); Chen et al. (2024); Huang et al. ; Liao et al. (2024); Hao et al. (2024); Zhu et al. (2024). Both approaches have their limitations. 1) The accumulation of low-rank updates requires instantiating a full-rank matrix and a deeply customized training strategy that periodically merges and restarts the low-rank components. This creates computing overhead in practice and can only achieve (if only) marginal computing and memory reduction. 2) Enforcing low-rank gradients reduces only the optimizer memory and adds additional computation that downgrades training throughput. Furthermore, the memory saving caused by gradient compression becomes negligible as the training batch size increases, as activations dominate the total memory cost. Recently SLTrain Han et al. (2024) revisited the notion of parameter efficiency in foundation model pre-training, by having both low-rank factors and an unstructured sparse matrix. SLTrain effectively reduces the total number of parameters without significantly hurting model performance. However, it still introduces computing overhead on top of full-rank training due to the necessary reconstruction of low-rank factors. We note that none of the above works has achieved superior efficiency of parameter, computing, and memory simultaneously without performance drop in both training and inference for foundation model pre-training.
@@
Contributions: In this paper, we rethink the fundamental architecture of LLMs and propose CoLA: Compute-Efficient Pre-Training of LLMs via Low-rank Activation, and its memory efficient implementation CoLA-M, to achieve all the desirable properties mentioned above. We summarize our contributions as follows:
  • We propose CoLA, a novel architecture to enforce explicit low-rank activations. LLMs use massive full-size MLP and linear layers. CoLA replaces them with auto-encoders. Each auto-encoder applies nonlinear activations between two low-rank factors, greatly reducing the parameter counts and computing FLOPS while performing on par with the full-rank pre-training.
@@
  • We extensively pre-train LLaMA (with 60M to 7B parameters) and BERT-large. CoLA reduces model size and computing FLOPs by $\bf 2\boldsymbol{\times}$, while maintaining on-par performance to its full-rank counterpart. At the system level, CoLA improves $\bf 1.86\boldsymbol{\times}$ training and $\bf 1.64\boldsymbol{\times}$ inference throughput. CoLA-M reduces total pre-training memory by $\bf 2/3$, while still manages to improve $\bf 1.3\boldsymbol{\times}$ training throughput over full-rank baselines.
A high-level comparison of CoLA(-M) with main baselines is provided in Table 1.
@@ 2 Related Work
Model Compression. Recent research on efficient LLM pre-training primarily focuses on memory savings. SLTrain Han et al. (2024) is the first method that reduces both trainable parameters and total parameters in LLM pre-training, without significantly hurting model performance. This also reduces memory usage for model, gradients, and optimizer states. However, the existence of its unstructured sparse matrix $\mathbf{S}$ requires reconstructing $\tilde{\mathbf{W}}=\mathbf{BA}+\mathbf{S}$, otherwise it will incur dense-sparse multiplications that are still memory costly (Fig. 3c). This causes additional computing than the full-rank baseline. LoRA/ReLoRA Hu et al. (2021); Lialin et al. (2023) reduces trainable parameters by freezing a full-rank $\mathbf{W}_{0}$ and training (at least in a later stage) only low-rank factors, potentially reducing memory needs. Yet, any compute savings are limited because the forward pass yields a larger compute than its full-rank counterpart, especially when the rank must stay relatively large in pre-training tasks. LoQT Loeschcke et al. (2024) further extends this formulation into quantized training. CoMERA Yang et al. (2024) achieves higher model compression and FLOPs reduction, yet its low-rank tensor operations are GPU unfriendly and can also cause a performance drop. Some works investigate pure structured sparsity or combined with low-rank factors Hu et al. (2024); Mozaffari et al. (2024), but still show a significant performance drop during the pre-training stage.
@@
Gradient Compression. GaLore Zhao et al. (2024) reduces memory by projecting gradients into a low-rank space, shrinking optimizer states below the typical 2×2$\times$2 × AdamW overhead Loshchilov (2017). However, it increases computation by adding up/down projections on top of already compute-heavy full-rank training. As shown in Fig. 1, its estimated FLOPs surpass full-rank training on the LLaMA-1B scale. Follow-up works Chen et al. (2024); Huang et al. ; Liao et al. (2024); Hao et al. (2024); Zhu et al. (2024) further explore low-rank gradient projection. While being promising, these methods are mostly orthogonal to our focus. Crucially, they are computing lower-bounded by the full-rank baseline. Our goal instead is to reduce computing cost to a fraction of full-rank LLM pre-training.
Activation Compression. CompAct Shamshoum et al. (2024) reduces memory of the computational graph using low-rank compression on saved activations, which introduces similar computing cost, yet underperforms GaLore. ESPACE Sakr and Khailany (2024) explores a very similar idea by projecting activations based on well-trained weight matrices, thus only applicable to the post-training stage. Crucially, the projections in both methods introduce additional computing costs on top of the full-rank baseline. And both of them do not change the fundamental structure of LLMs.
@@
This paper presents an architectural innovation that explicitly enforces low-rank activations by adopting the bottleneck-shaped auto-encoders as the building brick of the transformer architecture. This is conceptually different from the above model compression methods despite of some similarities in their formulations. Our approach is mostly orthogonal with gradient compression techniques, meaning that they could be combined to further boost efficiency.
@@ 3 CoLA for Efficient LLM Pre-Training
Figure 3: Comparison between different efficient pre-training frameworks. a) LoRA/ReLoRA Lialin et al. (2023) freezes a full-rank weight; b) GaLore Zhao et al. (2024) only reduces optimizer states by down and up projecting gradients; c) SLTrain Han et al. (2024) requires reconstruction of the low-rank and sparse matrices; d) CoLA (ours) is a pure low-rank architecture involving only rank $r$ weight matrices.
@@ 3.1 A Motivating Example
Many works have observed the low-rank structure of model activations in deep neural networks Cui et al. (2020); Huh et al. (2021). We also observe this phenomenon in LLMs, i.e. the effective rank of the activations is much smaller than their original dimensionality. To quantify this, we define the effective rank $r(\alpha)$ of activation as the minimal number of singular values needed to preserve an $\alpha$-fraction of the total spectral energy. Formally:
@@
3.2 Low-Rank Activation via Auto-Encoder
The above observation motivates us to ask one fundamental question: do we really need these full-size MLP and linear layers in LLMs? To eliminate the redundant activations, we propose to replace them with bottleneck-structured auto-encoders that naturally facilitate low-rank activations.
@@ 3.2 Low-Rank Activation via Auto-Encoder
We replace this MLP layer with an auto-encoder layer which consists low-rank matrices $\mathbf{A}\in\mathbb{R}^{r\times d_{\text{in}}}$ and $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ and a non-linear activation $\sigma$ in the middle. Rank $r<\min(d_{\text{in},\text{out}})$ is a design parameter that trades off between compute and performance. Formally, it can be written as:
$\mathbf{h^{\prime}}=\mathbf{B}\,\sigma(\mathbf{A}\mathbf{x}),$ (3)
The auto-encoder layer naturally enforces a low-rank activation in training, offering a principled approach to eliminate the redundancy observed in Fig. 2. We have the following remarks
  • The auto-encoder layer fundamentally differs from performing low-rank weight compression in an MLP layer. The latter performs lossy compression on model parameters but cannot eliminate the redundancy in activations.
  • The auto-encoder is not equivalent to using smaller feature dimensions in MLP layers, since $\mathbf{B}$ in the current layer cannot be merged with $\mathbf{A}$ in the next layer, due to the existence of various operations (e.g. residual connection) in the original dimension.
Since the low-rank property is widely observed regardless of whether $\mathbf{Wx}$ being followed by nonlinearity (see details in Appendix 11), we also uniformly adopt this auto-encoder structure to all projection layers in the transformer architecture. We empirically find that adding the original nonlinearity on top of Eq. (3) does not harm the performance, nor necessarily brings benefit (c.f. Appendix 10).
Figure 4: A decoder block in CoLA with LLaMA-like architecture (layer norms, rotary positional embeddings are omitted for simplicity). All MLP layers and projection layers in attention are replaced with auto-encoders. Modules painted in sketch are the re-computations during the backward step of CoLA-M (a memory efficient implementation of CoLA).
Fig. 4 shows the architecture of each transformer block when adopting CoLA into the LLaMA architecture. We highlight the fact that only the original linear layers and (if any) their follow-up non-linear transformation are modified to the CoLA formulation. Other computations such as the scaled-dot product of the self-attention, as well as residual connections and the element-wise product of LLaMA’s MLP layers, remain unchanged.
@@ 3.3 Computing Efficiency
We summarize the computational costs of these methods in Table 3 and observe that the costs of SLTrain and GaLore are lower bounded by full-rank training, while (Re)LoRA is lower bounded by CoLA when choosing the same rank. In contrast, CoLA reduces the computation from full-rank training when $r<0.62d$, assuming $d_{\text{ff}}\approx 2.5d$ in LLaMA-like architecture. The default rank choice is set to $r=\frac{1}{4}d$, leading to a reduction in compute to about half the full-rank training. We refer all details of compute analysis to Appendix B.
@@ 4 CoLA-M: A Memory-Efficient Implementation
In this section, we design and develop CoLA-M, a memory-efficient implementation to leverage CoLA’s structural advantage to achieve superior memory saving without sacrificing throughput.
@@ 4.1 Memory Breakdown in Pre-Training
Figure 6: Memory breakdown of pre-training LLaMA-1B on single GPU using different pre-training methods.
We assume a common notion that training modern transformers with Adam (or AdamW) involves four key memory components Zhao et al. (2024); Han et al. (2024): model parameters (1×1$\times$1 ×), gradients (1×1$\times$1 ×), optimizer states (2×2$\times$2 ×), and activations (14×1$\sim$ 4$\times$1 ∼ 4 ×). We focus on the scenario where the memory cost determined by the model size is not on the extreme limit of the GPU. We argue that this is rather realistic, since the model size and the minimum required tokens should scale up simultaneously during pre-training Kaplan et al. (2020); Hoffmann et al. (2022); Krajewski et al. (2024); Kumar et al. (2024). A tiny batch size on a single GPU would be impractical. Therefore, we analyze memory usage on a 40-GB A100 or a 94-GB H100 GPU with a fairly large sequence batch size. Fig. 5 & 6 show that activations dominate memory usage in this setup.